log3 64-log3 8/3+ log3 2=log3 4r
Answer by RAY100(1637)

log(3) 64 -log(3) 8/3 + log(3) 2 = log(3) 4R
log (3) [ { 64 *2} / (8/3) ] = log (3) 4R
log(3) (48) = log(3) 4R
48 = 4 R
R = 12,,,,,,,,ANSWER
,
checking
log 64/log3 -log(8/3)/log3 +log2//og3 = log 48/log 3
1.68 = 1.68 ,,,,,,ok